OpenStudy (anonymous):

3/5p-1/2=7/10 Help????

OpenStudy (anonymous):

3/5p=2.4 now you solve easily

OpenStudy (anonymous):

@vipul92 how do i slove it ?

OpenStudy (anonymous):

you can right 3/5p=(7/10)+(1/2) now solve it

OpenStudy (anonymous):

\[\frac{ 3 }{ 5p }- \frac{ 1 }{ 2 } = \frac{ 7 }{ 10 }\] find and lcd for 3/5p - 1/2 which is ???

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

10P cuz that variable \[\frac{ 6 }{ 10p } - \frac{ 5p }{ 10p } = \frac{ 7p }{ 10p }\] combine like terms